The length of a rectangular rug is 6 feet less than the width. If the aera of the rug is 72 ft^2, find the dimensions.
Answer by nerdybill(7384)

The length of a rectangular rug is 6 feet less than the width. If the aera of the rug is 72 ft^2, find the dimensions.
.
Let w = width
then
w-6 = length
.
w(w-6) = 72
w^2 - 6w = 72
w^2 - 6w - 72 = 0
(w+6)(w-12) = 0
w = {-6, 12}
throw out the negative solution leaving:
w = 12 ft (width)
length:
w-6 = 12-6 = 6 ft (length)